# Next-to-leading order analysis of $J/\psi+\gamma$ production in photon-photon collisions at CEPC

### v1

@@ Abstract
We systematically investigate the production of $J/\psi+\gamma$ in $\gamma\gamma$ collisions at next-to-leading order in $\alpha_{s}$ within nonrelativistic QCD (NRQCD) factorization. Calculations for CEPC energy region show the resolved photon contribution is negligible, while the direct photon process dominates, yielding substantial annual $J/\psi$ yields. Significant modifications to $J/\psi$ polarization parameters emerge from color-octet mechanisms, and different NRQCD long distance matrix elements (LDMEs) further yield distinct polarization patterns. Furthermore, the polarization predictions are highly sensitive to the $^{3}P_{J}^{[8]}$ LDME, while being insensitive to the $^{1}S_{0}^{[8]}$ and $^{3}S_{1}^{[8]}$ LDMEs. Leveraging the cleaner environment of $e^{+}e^{-}$ collisions versus hadronic processes, the production of $J/\psi$ associated with a photon in $\gamma\gamma$ collisions provides a high-precision platform to test LDMEs universality and resolve longstanding $J/\psi$ polarization puzzles.
@@ I Introduction
In this article, we propose utilizing the process of $J/\psi$ production associated with an isolated photon via two-photon collisions to further investigate the polarization of $J/\psi$. At $e^{+}e^{-}$ colliders, such collisions are mediated by nearly on-shell virtual photons from incoming leptons, offering a cleaner environment than hadroproduction due to suppressed strong-interaction backgrounds. For the specific process of $\gamma\gamma\to J/\psi+\gamma$, the direct channel dominates over single- and double-resolved contributions. This suppression stems from CO LDMEs and the Parton Distribution Distributions (PDFs) of gluon/quark in a resolved photon. In contrast, inclusive $J/\psi$ production in $\gamma\gamma$ collisions is dominated by single-resolved processes, introducing theoretical uncertainties from the poorly constrained gluon/quark PDFs in photon. As will be demonstrated, the $J/\psi$ polarization prediction reveals relevance of the $^{3}P_{J}^{[8]}$ state with negligible $^{1}S_{0}^{[8]}$ contributions, while $^{3}S_{1}^{[8]}$ is strongly suppressed by $\mathcal{O}(\alpha_{s})$ higher-order effects. These features establish the $J/\psi+\gamma$ production in photon-photon collision as a clean probe for investigating $J/\psi$ polarization. Considering its designed high luminosity and collision energy at CEPC, which offers large potential for the production of a statistically significant sample of $J/\psi+\gamma$ events via photon-photon collisions, this study will focus on the CEPC CEPCStudyGroup:2018rmc ; CEPCStudyGroup:2018ghi as the proposed platform.
@@
The process of $J/\psi$ production accompanied by a photon has been extensively studied across various collider environments Drees:1991ig ; Doncheski:1993dj ; Kim:1994bm ; Roy:1994vb ; Mirkes:1994jr ; Kim:1996bb ; Mehen:1996vx ; Cacciari:1996zu ; Ma:1997bi ; Japaridze:1998ss ; Mathews:1999ye ; Kniehl:2002wd ; Klasen:2004az ; Kniehl:2006qq ; Li:2008ym ; Lansberg:2009db ; Li:2014ava ; Alimov:2024pqt ; Keung:1983ac ; Bodwin:2013gca ; Bodwin:2014bpa ; Huang:2014cxa ; Bodwin:2016edd ; Bodwin:2017wdu ; Bodwin:2017pzj ; Li:2009ki ; Sang:2009jc ; Braguta:2010mf ; Patel:2011zty ; Fan:2012vw ; Fan:2012dy ; Chao:2013cca ; Chen:2013mjb ; Li:2013nna ; Chen:2013itc ; Wang:2013ywc ; Sun:2014kva ; Xu:2014zra ; Chen:2017pyi ; Brambilla:2017kgw ; Chung:2019ota ; Li:2019ncs ; Brambilla:2020xod ; Yu:2020tri ; Sang:2020fql ; Liao:2021ifc ; Wang:2023ssg . Researches at hadronic colliders demonstrate significant discrepancies between NRQCD and CS predictions for the $pp\to J/\psi+\gamma+X$ process Kim:1996bb ; Cacciari:1996zu ; Li:2014ava , particularly in polarization distributions Li:2014ava . Studies at electron-proton colliders (e.g., HERA) have also revealed the crucial role of the CO mechanism in the $ep\to J/\psi+\gamma+X$ process Kniehl:2006qq . Furthermore, the production of $J/\psi$ accompanied by a photon from $Z$ or Higgs boson decays has been systematically investigated Keung:1983ac ; Bodwin:2013gca ; Bodwin:2014bpa ; Huang:2014cxa ; Bodwin:2016edd ; Bodwin:2017wdu ; Bodwin:2017pzj . In electron-positron annihilation processes, heavy quarkonium production with an associated photon have also been wildly studied Li:2009ki ; Sang:2009jc ; Braguta:2010mf ; Patel:2011zty ; Fan:2012vw ; Fan:2012dy ; Chao:2013cca ; Chen:2013mjb ; Li:2013nna ; Chen:2013itc ; Wang:2013ywc ; Sun:2014kva ; Xu:2014zra ; Chen:2017pyi ; Brambilla:2017kgw ; Chung:2019ota ; Li:2019ncs ; Brambilla:2020xod ; Yu:2020tri ; Sang:2020fql ; Liao:2021ifc ; Wang:2023ssg . Building upon this foundation, the complete NLO calculations within the NRQCD framework for the $\gamma\gamma\to J/\psi+\gamma$ process in photon-photon collisions, with particular focus on $J/\psi$ polarization, will thereby extend the theoretical research scope of $J/\psi$-photon associated production processes.
@@ II Calculation Formalism
We adopt three representative sets of LDMEs in our calculations, each providing consistent predictions for $J/\psi$ hadroproductions. Their values are listed in Table I, with the LDMEs of Refs. Butenschoen:2010rq , Chao:2012iv , and Zhang:2014ybe labeled as set 1, set 2, and set 3, respectively.
@@
Figure 1 displays the differential cross sections for direct, single-, and double-resolved processes. The direct channel dominates by at least two orders of magnitude, demonstrating that $\gamma\gamma\to J/\psi+\gamma$ proceeds predominantly through direct production. This dominance arises from three key factors: 1) the direct process $\gamma\gamma\to c\bar{c}[^{3}S_{1}^{[1]}]+\gamma$ occurs at leading order via color-singlet mechanisms; 2) the single-resolved production $\gamma+g\to c\bar{c}[^{3}S_{1}^{[8]}]+\gamma$ suffers strong suppression from $\langle\mathcal{O}^{J/\psi}(^{3}S_{1}^{[8]})\rangle/\langle\mathcal{O}^{J/\psi}(^{3}S_{1}^{[1]})\rangle$; 3) the double-resolved CS process $gg\to c\bar{c}[^{3}S_{1}^{[1]}]+\gamma$, mediated by gluon fusion, is suppressed by the small gluon PDF of the photon. In light of these suppression mechanisms, we exclusively analyze the direct channel in subsequent calculations.
@@
For the virtual corrections to the process $\gamma+\gamma\to c\bar{c}[^{3}S_{1}^{[1]}]+\gamma$, there are 78 Feynman diagrams in total, comprising 48 one-loop diagrams and 30 counter-term diagrams, as shown representatively in Figure 2. To isolate the ultraviolet (UV) and infrared (IR) singularities from the one-loop diagrams, dimensional regularization with $D=4-2\epsilon$ is employed. The on-mass-shell (OS) scheme is adopted for the renormalization constants of the quark mass $Z_{\textrm{m}}$ and the quark field $Z_{2}$, given by
@@
For the process $\gamma+\gamma\to c\bar{c}[^{1}S_{0}^{[8]},^{3}P_{J}^{[8]}]+\gamma+g$, there are 24 Feynman diagrams, as shown representatively in Figure 3. The $^{1}S_{0}^{[8]}$ channel is divergence free, permitting direct computation. However, in the $^{3}P_{J}^{[8]}$ channel, the emitted gluon can be soft, leading to IR singularities. These singularities are absorbed through the renormalization of the corresponding NRQCD matrix elements $\langle\mathcal{O}^{{}^{3}P_{J}^{[8]}}(^{3}S_{1}^{[1]})\rangle$ . [^1] Owing the absence of $^{3}S_{1}^{[8]}-$related processes up to $\mathcal{O}(\alpha^{3}\alpha_{s})$, only $\langle\mathcal{O}^{{}^{3}P_{J}^{[8]}}(^{3}S_{1}^{[1]})\rangle$ requires renormalization. To derive $\hat{\sigma}_{{}^{3}P_{J}^{[8]}}$, we start by decomposing the cross section $\sigma(^{3}P_{J}^{[8]})$ into two components,
@@
$\hat{\sigma}_{\textrm{S}}$ denotes the soft terms, which can be expressed by means of the LO SDC $\hat{\sigma}^{\mathrm{LO}}_{{}^{3}S_{1}^{[1]}}$ for the partonic process $\gamma\gamma\to c\bar{c}[{}^{3}S_{1}^{[1]}]+\gamma$. $\hat{\sigma}_{\textrm{H}}$ represents the hard contributions, which are IR finite and can be computed numerically using standard Monte Carlo integration techniques.
@@
$\hat{\sigma}_{{}^{3}P_{J}^{[8]}}=\hat{\sigma}_{\textrm{H}}-\frac{4\alpha_{s}}{3\pi m^{2}_{c}}u_{\epsilon}\hat{\sigma}^{\textrm{LO}}_{{}^{3}S_{1}^{[1]}}.$ (6)
The factor $u_{\epsilon}$ is finite, and its analytic expression can be found in Reference Zhang:2014coi .
In our calculations, we employ our Mathematica-Fortran package, which implements FeynArts Hahn:2000kx , FeynCalc Mertig:1990an , FIRE Smirnov:2008iw , and Apart Feng:2012iq . This package has been used for several heavy-quarkonium processes in $\gamma\gamma$ and $\gamma p$ collisions Sun:2017nly ; Sun:2017wxk ; Sun:2015hhv . Additionally, we utilize the independent Feynman Diagram Calculation (FDC) package Wang:2004du to compute all relevant processes, obtaining consistent numerical results.
@@ III Phenomenological results
The integrated cross sections of $J/\psi+\gamma$ production in $\gamma\gamma$ collisions, computed for three LDME sets, are presented in Table 2. “CS” denotes the contribution of $c\bar{c}[^{3}S_{1}^{[1]}]$, while “NRQCD” represents the complete calculation including both CS and CO ($^{1}S_{0}^{[8]}$, $^{3}P_{J}^{[8]}$) channels. The tabulated data reveal that NLO corrections to the $^{3}S_{1}^{[1]}$ process significantly suppress the LO results, with a reduction magnitude approaching 50%. Contributions from CO states exhibit slight impact on the integrated cross section. Assuming the nominal integrated luminosity of 300 fb-1 CEPCStudyGroup:2018rmc ; CEPCStudyGroup:2018ghi and considering the dileptonic branching fraction $\mathcal{B}(J/\psi\to\ell^{+}\ell^{-})\approx 10\%$, CEPC is expected to accumulate $\mathcal{O}(10^{3})$ observable $J/\psi$ events per year.
@@
The $p_{t}$ distributions in Figure 4 compare NRQCD predictions with CS results, demonstrating a rapid decrease in differential cross section $d\sigma/dp_{t}$ with increasing transverse momentum. Slight CO contributions cause near-overlap between CS and full NRQCD curves.
@@
Figure 5 reveals close agreement between CSM and full NRQCD polarization predictions for LDME sets 1 and 2. In contrast, set 3 exhibits a significant discrepancy: while CSM yields pronounced transverse polarization, NRQCD indicates nearly unpolarized production ($\lambda_{\theta}\sim 0$). In consideration of the consistent yield predictions between CS and NRQCD frameworks, the polarization divergence in set 3 necessitates investigation into its physical origin.
@@
To address this, we decompose the differential cross sections into transverse (T) and longitudinal (L) components, as compared in Figure 6. For sets 1 and 2, CO contributions marginally perturb both the transverse and longitudinal CS components. For set 3, however, CO terms substantially suppress the “T” parts while enhancing “L” parts, thereby transforming the initially transverse-polarized state into near unpolarization. Equation 6 reveals that $\hat{\sigma}_{{}^{3}P_{J}^{[8]}}$ contains a negative term proportional to $\hat{\sigma}_{{}^{3}S_{1}^{[1]}}^{\textrm{LO}}$. Owing to the transverse dominance in $\hat{\sigma}_{{}^{3}S_{1}^{[1]}}^{\textrm{LO}}$, $\hat{\sigma}_{{}^{3}P_{J}^{[8]}}$ therefore develops negative transverse and positive longitudinal components. For LDME sets 1 and 2, the small ratio $R=\frac{\langle\mathcal{O}^{J/\psi}(^{3}P_{J}^{[8]})\rangle}{\langle\mathcal{O}^{J/\psi}(^{3}S_{1}^{[1]})\rangle}$ leads to modest polarization modifications. In contrast, set 3 exhibits a relatively large ratio $R$, where the $^{3}P_{J}^{[8]}$ contribution significantly suppresses the transverse component while enhancing the longitudinal one, subsequently drastically altering the CS polarization pattern. Crucially, since the $^{1}S_{0}^{[8]}$ contribution is negligible and unpolarized, and $^{3}S_{1}^{[8]}$ states are absent at current $\alpha_{s}$ order, $J/\psi$ polarization in $\gamma\gamma\to J/\psi+\gamma+X$ is solely sensitive to $R$. The persistent challenges in describing $J/\psi$ polarization largely stem from the difficulties in determining the LDMEs. A key advantage of the process $\gamma\gamma\to J/\psi+\gamma+X$ over hadroproduction is its selective sensitivity: while the latter receives essential contributions from all three CO states ($^{1}S_{0}^{[8]}$, $^{3}S_{1}^{[8]}$, and $^{3}P_{J}^{[8]}$), polarization in the former depends almost solely on the $^{3}P_{J}^{[8]}$ LDME, consequently providing a cleaner probe for constraining this specific matrix element and improving the precision of NRQCD predictions.
@@
Finally, Figure 7 compares polarization parameters predicted by the three LDME sets, incorporating their respective uncertainties. The uncertainty bands reveal distinct polarization patterns:

### v2

@@ Abstract
We systematically investigate the production of $J/\psi+\gamma$ in $\gamma\gamma$ collisions within nonrelativistic QCD (NRQCD) factorization, with the direct-photon channel calculated specifically up to the next-to-leading order in $\alpha_{s}$. Calculations for CEPC energy region show the resolved photon contribution is negligible, while the direct photon process dominates, yielding substantial annual $J/\psi$ yields. Significant modifications to $J/\psi$ polarization parameters emerge from color-octet mechanisms, and different NRQCD long distance matrix elements (LDMEs) further yield distinct polarization patterns. Furthermore, the polarization predictions are highly sensitive to the $^{3}P_{J}^{[8]}$ LDME, while being insensitive to the $^{1}S_{0}^{[8]}$ and $^{3}S_{1}^{[8]}$ LDMEs. Leveraging the cleaner environment of $e^{+}e^{-}$ collisions versus hadronic processes, the production of $J/\psi$ associated with a photon in $\gamma\gamma$ collisions provides a high-precision platform to test LDMEs universality and resolve longstanding $J/\psi$ polarization puzzles.
@@ I Introduction
In this article, we propose utilizing the process of $J/\psi$ production associated with an isolated photon via two-photon collisions to further investigate the polarization of $J/\psi$. At $e^{+}e^{-}$ colliders, such collisions are mediated by nearly on-shell virtual photons from incoming leptons, offering a cleaner environment than hadroproduction due to suppressed strong-interaction backgrounds. For the specific process of $\gamma\gamma\to J/\psi+\gamma$, the direct channel dominates over single- and double-resolved contributions. This suppression stems from CO LDMEs and the Parton Distribution Functions (PDFs) of gluon/quark in a resolved photon. In contrast, inclusive $J/\psi$ production in $\gamma\gamma$ collisions is dominated by single-resolved processes, introducing theoretical uncertainties from the poorly constrained gluon/quark PDFs in photon. As will be demonstrated, the $J/\psi$ polarization prediction reveals relevance of the $^{3}P_{J}^{[8]}$ state with negligible $^{1}S_{0}^{[8]}$ contributions, while $^{3}S_{1}^{[8]}$ is strongly suppressed by $\mathcal{O}(\alpha_{s})$ higher-order effects. These features establish the $J/\psi+\gamma$ production in photon-photon collision as a clean probe for investigating $J/\psi$ polarization. Considering its designed high luminosity and collision energy at CEPC, which offers large potential for the production of a statistically significant sample of $J/\psi+\gamma$ events via photon-photon collisions, this study will focus on the CEPC CEPCStudyGroup:2018rmc ; CEPCStudyGroup:2018ghi as the proposed platform.
@@
The process of $J/\psi$ production accompanied by a photon has been extensively studied across various collider environments Drees:1991ig ; Doncheski:1993dj ; Kim:1994bm ; Roy:1994vb ; Mirkes:1994jr ; Kim:1996bb ; Mehen:1996vx ; Cacciari:1996zu ; Ma:1997bi ; Japaridze:1998ss ; Mathews:1999ye ; Kniehl:2002wd ; Klasen:2004az ; Kniehl:2006qq ; Li:2008ym ; Lansberg:2009db ; Li:2014ava ; Alimov:2024pqt ; Keung:1983ac ; Bodwin:2013gca ; Bodwin:2014bpa ; Huang:2014cxa ; Bodwin:2016edd ; Bodwin:2017wdu ; Bodwin:2017pzj ; Li:2009ki ; Sang:2009jc ; Braguta:2010mf ; Patel:2011zty ; Fan:2012vw ; Fan:2012dy ; Chao:2013cca ; Chen:2013mjb ; Li:2013nna ; Chen:2013itc ; Wang:2013ywc ; Sun:2014kva ; Xu:2014zra ; Chen:2017pyi ; Brambilla:2017kgw ; Chung:2019ota ; Li:2019ncs ; Brambilla:2020xod ; Yu:2020tri ; Sang:2020fql ; Liao:2021ifc ; Wang:2023ssg . Researches at hadronic colliders demonstrate significant discrepancies between NRQCD and CS predictions for the $pp\to J/\psi+\gamma+X$ process Kim:1996bb ; Cacciari:1996zu ; Li:2014ava , particularly in polarization distributions Li:2014ava . Studies at electron-proton colliders (e.g., HERA) have also revealed the crucial role of the CO mechanism in the $ep\to J/\psi+\gamma+X$ process Kniehl:2006qq . Furthermore, the production of $J/\psi$ accompanied by a photon from $Z$ or Higgs boson decays has been systematically investigated Keung:1983ac ; Bodwin:2013gca ; Bodwin:2014bpa ; Huang:2014cxa ; Bodwin:2016edd ; Bodwin:2017wdu ; Bodwin:2017pzj . In electron-positron annihilation processes, heavy quarkonium production with an associated photon have also been widely studied Li:2009ki ; Sang:2009jc ; Braguta:2010mf ; Patel:2011zty ; Fan:2012vw ; Fan:2012dy ; Chao:2013cca ; Chen:2013mjb ; Li:2013nna ; Chen:2013itc ; Wang:2013ywc ; Sun:2014kva ; Xu:2014zra ; Chen:2017pyi ; Brambilla:2017kgw ; Chung:2019ota ; Li:2019ncs ; Brambilla:2020xod ; Yu:2020tri ; Sang:2020fql ; Liao:2021ifc ; Wang:2023ssg . In 2005, Klasen et al. performed the first analytical NLO QCD calculation for $\gamma\gamma\to J/\psi+\gamma$ Klasen:2004az , highlighting its promising detectability at the $e^{+}e^{-}$ linear collider TESLA. In this article, we extend the theoretical scope by presenting the NRQCD-based NLO calculation of $J/\psi$ polarization in this process at the $e^{+}e^{-}$ circular collider CEPC, thereby advancing the understanding of quarkonium-photon associated production.
@@ II Calculation Formalism
set 4 Brambilla:2024iqg $1.16\pm 0.068$ $0.2489\pm 0.34$ $1.050\pm 0.121$ $1.879\pm 0.261$
We employ four representative LDME sets (labeled set 1–4), each providing a consistent description of $J/\psi$ hadroproduction data. Their values, corresponding respectively to the fits in Refs. Butenschoen:2010rq ; Chao:2012iv ; Zhang:2014ybe ; Brambilla:2024iqg , are listed in Table I.
@@
Fig. 1 displays the differential cross sections for direct, single-, and double-resolved processes. The direct channel dominates by at least two orders of magnitude, demonstrating that $\gamma\gamma\to J/\psi+\gamma$ proceeds predominantly through direct production. This dominance arises from three key factors: 1) the direct process $\gamma\gamma\to c\bar{c}[^{3}S_{1}^{[1]}]+\gamma$ occurs at leading order via color-singlet mechanisms; 2) the single-resolved production $\gamma+g\to c\bar{c}[^{3}S_{1}^{[8]}]+\gamma$ suffers strong suppression from $\langle\mathcal{O}^{J/\psi}(^{3}S_{1}^{[8]})\rangle/\langle\mathcal{O}^{J/\psi}(^{3}S_{1}^{[1]})\rangle$; 3) the double-resolved CS process $gg\to c\bar{c}[^{3}S_{1}^{[1]}]+\gamma$, mediated by gluon fusion, is suppressed by the small gluon PDF of the photon. In light of these suppression mechanisms, we exclusively analyze the direct channel in subsequent calculations.
@@
For the virtual corrections to the process $\gamma+\gamma\to c\bar{c}[^{3}S_{1}^{[1]}]+\gamma$, there are 78 Feynman diagrams in total, comprising 48 one-loop diagrams and 30 counter-term diagrams, as shown representatively in Fig. 2. To isolate the ultraviolet (UV) and infrared (IR) singularities from the one-loop diagrams, dimensional regularization with $D=4-2\epsilon$ is employed. The on-mass-shell (OS) scheme is adopted for the renormalization constants of the quark mass $Z_{\textrm{m}}$ and the quark field $Z_{2}$, given by
@@
For the process $\gamma+\gamma\to c\bar{c}[^{1}S_{0}^{[8]},^{3}P_{J}^{[8]}]+\gamma+g$, there are 24 Feynman diagrams, as shown representatively in Fig. 3. The $^{1}S_{0}^{[8]}$ channel is divergence free, permitting direct computation. However, in the $^{3}P_{J}^{[8]}$ channel, the emitted gluon can be soft, leading to IR singularities. These singularities are absorbed through the renormalization of the corresponding NRQCD matrix elements $\langle\mathcal{O}^{{}^{3}P_{J}^{[8]}}(^{3}S_{1}^{[1]})\rangle$ . [^1] Owing the absence of $^{3}S_{1}^{[8]}-$related processes up to $\mathcal{O}(\alpha^{3}\alpha_{s})$, only $\langle\mathcal{O}^{{}^{3}P_{J}^{[8]}}(^{3}S_{1}^{[1]})\rangle$ requires renormalization. To derive $\hat{\sigma}_{{}^{3}P_{J}^{[8]}}$, we start by decomposing the cross section $\sigma(^{3}P_{J}^{[8]})$ into two components,
@@
The small cutoff parameter $\delta_{s}$ is introduced to separate the phase space into soft ($\hat{\sigma}_{\text{S}}$) and hard ($\hat{\sigma}_{\text{H}}$) regions. The soft part $\hat{\sigma}_{\text{S}}$ can be expressed analytically in terms of the LO SDC $\hat{\sigma}^{\mathrm{LO}}_{{}^{3}S_{1}^{[1]}}$ for the partonic process $\gamma\gamma\to c\bar{c}[{}^{3}S_{1}^{[1]}]+\gamma$. The hard part $\hat{\sigma}_{\text{H}}$ is infrared-finite and can be evaluated numerically using standard Monte Carlo integration techniques.
@@
$\hat{\sigma}_{{}^{3}P_{J}^{[8]}}=\hat{\sigma}_{\textrm{H}}-\frac{4\alpha_{s}}{3\pi m^{2}_{c}}u_{\epsilon}\hat{\sigma}^{\textrm{LO}}_{{}^{3}S_{1}^{[1]}},$ (6)
with Zhang:2014coi
$u_{\epsilon}=\frac{p_{0}}{|\bf{p}|}\textrm{ln}\left(\frac{p_{0}+|\bf{p}|}{p_{0}-|\bf{p}|}\right)+\textrm{ln}\left(\frac{\mu_{\Lambda}^{2}}{s\delta_{s}^{2}}\right)-2+2\textrm{ln}(2),$ (7)
where $p_{0}$ and $\mathbf{p}$ denote the energy and three-momentum of the $J/\psi$, respectively. $\mu_{\Lambda}$ denotes the scale arising from the renormalization of the LDME and is set to $\mu_{\Lambda}=m_{c}$.
In our calculations, we employ our Mathematica-Fortran package, which implements FeynArts Hahn:2000kx , FeynCalc Mertig:1990an , FIRE Smirnov:2008iw , and Apart Feng:2012iq . This package has been used for several heavy-quarkonium processes in $\gamma\gamma$ and $\gamma p$ collisions Sun:2017nly ; Sun:2017wxk ; Sun:2015hhv . Additionally, we utilize the independent Feynman Diagram Calculation (FDC) package Wang:2004du to compute all relevant processes, obtaining consistent numerical results. By coherently superimposing the bremsstrahlung Frixione:1993yw and beamstrahlung Chen:1993dba photon spectra as described in Ref. Klasen:2004az , we have performed a comparison with the results therein. We observe a discrepancy: our LO and NLO cross sections are systematically lower by $15\%$$35\%$ compared to Fig. 8(a) of Ref. Klasen:2004az . Nevertheless, the $K$-factors we obtain show trends similar to those in its Fig. 9(a). As a crosscheck of our computational framework, we reproduced the LO and NLO cross sections for the direct process $\gamma\gamma\to c\bar{c}[^{3}S_{1}^{[1]}]+\gamma$ from Fig. 3 of Ref. Yedelkina:2023cbu under identical kinematic cuts.
@@ III Phenomenological results
In this section, we present our numerical results of the process $\gamma\gamma\to J/\psi+\gamma$, which exclusively consider the direct photon production at NLO accuracy and exclude the resolved photon contributions.
@@
set 4 $0.197$ $0.095$ $0.091$
The integrated cross sections of $J/\psi+\gamma$ production in $\gamma\gamma$ collisions, computed for four LDME sets, are presented in Table 2. “CS” denotes the contribution of $c\bar{c}[^{3}S_{1}^{[1]}]$, while “NRQCD” represents the complete calculation including both CS and CO ($^{1}S_{0}^{[8]}$, $^{3}P_{J}^{[8]}$) channels. The tabulated data reveal that NLO corrections to the $^{3}S_{1}^{[1]}$ process significantly suppress the LO results, with a reduction magnitude approaching 50%. Contributions from CO states exhibit slight impact on the integrated cross section. Assuming the nominal integrated luminosity of 300 fb-1 CEPCStudyGroup:2018rmc ; CEPCStudyGroup:2018ghi and considering the dileptonic branching fraction $\mathcal{B}(J/\psi\to\ell^{+}\ell^{-})\approx 10\%$, CEPC is expected to accumulate $\mathcal{O}(10^{3})$ observable $J/\psi$ events per year.
@@
The $p_{t}$ distributions in Fig. 4 compare NRQCD predictions with CS results, demonstrating a rapid decrease in differential cross section $d\sigma/dp_{t}$ with increasing transverse momentum. Slight CO contributions cause near-overlap between CS and full NRQCD curves.
@@
Fig. 5 reveals close agreement between CSM and full NRQCD polarization predictions for LDME sets 1 and 2. By contrast, a significant discrepancy emerges for sets 3 and 4: while the CSM yields pronounced transverse polarization, the full NRQCD calculation gives nearly unpolarized production for set 3 and only weakly transverse polarization for set 4. In consideration of the consistent yield predictions between CS and NRQCD frameworks, the polarization divergence in sets 3 and 4 necessitates investigation into its physical origin.
@@
To address this, we decompose the differential cross sections into transverse (T) and longitudinal (L) components, as compared in Fig. 6. For sets 1 and 2, CO contributions marginally perturb both the transverse and longitudinal CS components. For sets 3 and 4, however, CO terms substantially suppress the “T” parts while enhancing “L” parts, thereby transforming the initially transverse-polarized state into near unpolarization. Equation 6 reveals that $\hat{\sigma}_{{}^{3}P_{J}^{[8]}}$ contains a negative term proportional to $\hat{\sigma}_{{}^{3}S_{1}^{[1]}}^{\textrm{LO}}$. Owing to the transverse dominance in $\hat{\sigma}_{{}^{3}S_{1}^{[1]}}^{\textrm{LO}}$, $\hat{\sigma}_{{}^{3}P_{J}^{[8]}}$ therefore develops negative transverse and positive longitudinal components. For LDME sets 1 and 2, the small ratio $R=\frac{\langle\mathcal{O}^{J/\psi}(^{3}P_{J}^{[8]})\rangle}{\langle\mathcal{O}^{J/\psi}(^{3}S_{1}^{[1]})\rangle}$ leads to modest polarization modifications. In contrast, sets 3 and 4 exhibit a relatively large ratio $R$, where the $^{3}P_{J}^{[8]}$ contribution significantly suppresses the transverse component while enhancing the longitudinal one, subsequently drastically altering the CS polarization pattern. Crucially, since the $^{1}S_{0}^{[8]}$ contribution is negligible and unpolarized, and $^{3}S_{1}^{[8]}$ states are absent at current $\alpha_{s}$ order, $J/\psi$ polarization in $\gamma\gamma\to J/\psi+\gamma+X$ is solely sensitive to $R$. The persistent challenges in describing $J/\psi$ polarization largely stem from the difficulties in determining the LDMEs. A key advantage of the process $\gamma\gamma\to J/\psi+\gamma+X$ over hadroproduction is its selective sensitivity: while the latter receives essential contributions from all three CO states ($^{1}S_{0}^{[8]}$, $^{3}S_{1}^{[8]}$, and $^{3}P_{J}^{[8]}$), polarization in the former depends almost solely on the $^{3}P_{J}^{[8]}$ LDME, consequently providing a cleaner probe for constraining this specific matrix element and improving the precision of NRQCD predictions.
@@
Finally, Fig. 7 compares polarization parameters predicted by the four LDME sets, incorporating their respective uncertainties. The uncertainty bands reveal distinct polarization patterns:
@@
  • set 4 produces near-unpolarized to slightly transverse polarization.